34.1 If a particular atom has a mass of 6.644 x 1026 kg and has a total of 20 electrons when neutral (uncharged)what element is it (use a periodic table to answer this question)?
Answer #1

Since given that element is neutral, So there will be only one element which will have 20 electrons.

from periodic table we can see that

Calcium ( Ca) has 20 electrons and 20 protons and 20 neutrons.

And mass of calcium = 40 amu = 40*1.67*10^-27 kg = 6.68*10^-26 kg

So this element will be Calcium (Ca)

Please Upvote.

Comment below if you have any query.